Hi! I need to calculate the discriminant and state how many real solutions each equation has.
4x^2-6x+4=0
and
4x^2-4x+1=0
Thanks for youe help in advance!

Answer by stanbon(75887)

calculate the discriminant and state how many real solutions each equation has.
4x^2-6x+4=0
a=4, b=-6, c=4
Discriminant= b^2-4ac = 6^2-4(4)(4)=36-64<0
Since it is <0 there are two complex roots, no Real Number solutions.
--------------------------------
and
4x^2-4x+1=0
a=4, b=-4, c=1
Discriminant= b^2-4ac = 16-4*4*1=0
Since it is 0 there is one Real Number solution with multiplicity two.
--------------------
Cheers,
Stan H.